A man leaves the garrage in his house and drives to a neighbouring town which is twenty kilometres away from his house on sight-seeing. He returns home to his garrage two hours after. What is his average velocity from home in km/h
1
Expert's answer
2015-05-15T03:05:28-0400
Solution:
average velocity = total distancetraveled / total time taken.
total time taken = 2 hours
total distance traveled =( distance from his house to a neighbouring town)+( distance from a neighbouring town to his house) =
20+20=40 km.
So average velocity = 40/2=20km/h